OpenStudy (anonymous):

what is the equation of a line tangent to the curve x^3+y^3=6xy at point (3,3) using implicit differentiation

jimthompson5910 (jim_thompson5910):

x^3+y^3=6xy 3x^2+3y^2*y'=6y+6xy' 3y^2*y'-6xy'=6y-3x^2 y'(3y^2-6x)=6y-3x^2 y'=(6y-3x^2)/(3y^2-6x) y'=(3(2y-x^2))/(3(y^2-2x)) y'=(2y-x^2))/(y^2-2x) y'=(2(3)-(3)^2)/((3)^2-2(3)) y'=-3/3 y' = -1 So the slope of the tangent line is -1 y = mx+b 3 = (-1)(3)+b 3 = -3 + b 6 = b b = 6 So the equation of the tangent line is y = -x+6

jimthompson5910 (jim_thompson5910):

alexray19 you forgot to derive 6xy to get 6y+6xy'

OpenStudy (anonymous):

Oh you're right, thanks :)

OpenStudy (anonymous):

thanks, having to refresh on my calc

OpenStudy (anonymous):

A solution using Mathematica, including a solution plot and comments is attached.